# Thunder-NUBench: A Benchmark for LLMs’ Sentence-Level Negation Understanding

### v3

@@ Abstract
Negation is a fundamental linguistic phenomenon that poses ongoing challenges for Large Language Models (LLMs), particularly in tasks requiring deep semantic understanding. Current benchmarks often treat negation as a minor detail within broader tasks, such as natural language inference. Consequently, there is a lack of benchmarks specifically designed to evaluate comprehension of negation. In this work, we introduce Thunder-NUBench — a novel benchmark explicitly created to assess sentence-level understanding of negation in LLMs. Thunder-NUBench goes beyond merely identifying surface-level cues by contrasting standard negation with structurally diverse alternatives, such as local negation, contradiction, and paraphrase. This benchmark includes manually curated sentence-negation pairs and a multiple-choice dataset, allowing for a comprehensive evaluation of models’ understanding of negation.
[…]
Thunder-NUBench: A Benchmark for
LLMs’ Sentence-Level Negation Understanding
Yeonkyoung So, Gyuseong Lee, Sungmok Jung, Joonhak Lee, JiA Kang, Sangho Kim, Jaejin Lee Graduate School of Data Science, Seoul National University {kathy1028,ksnannaya,tjdahrwjd,hmjelee,jia6776,ksh4931,jaejin}@snu.ac.kr
@@ 1 Introduction
Negation is a fundamental and universal phenomenon found in languages worldwide. It is closely associated with various human communicative abilities, such as denial, contradiction, deception, misrepresentation, and irony. Although affirmative statements are more common, negation still plays a significant role in language; approximately 25% of sentences in English texts contain some form of negation (sarabi2016understanding; hossain2020analysis; sep-negation). This prevalence and its impact on meaning make accurate interpretation of negation crucial for several natural language processing (NLP) tasks, including sentiment analysis, question answering, knowledge base completion, and natural language inference (NLI) (khandelwal2020negbert; hosseini2021understanding; singh2023nlms). Recent studies have shown that effectively managing negation is important even for multimodal language models (quantmeyer2024and; alhamoud2025vision; park2025know).
Meanwhile, negation poses significant challenges for both humans and language models. Research shows that people often find it more difficult to process and comprehend negated statements compared to affirmative ones (wales1969so; sarabi2016understanding). Similarly, many studies indicate that pretrained language models (PLMs) struggle to interpret negation accurately. For example, models like BERT (devlin2019bert) and even large language models (LLMs) such as GPT-3 (brown2020language) often have difficulty distinguishing between negated and affirmative statements. These models tend to rely on superficial cues, which can result in incorrect outputs when negation is involved (kassner2020negated; hossain2022analysis; truong2023language).
Figure 1: An example of Thunder-NUBench multiple-choice evaluation task, where the underlined text indicates the main verb phrase of each sentence, and the red text marks the negated part.
Despite its significance, there is a notable lack of dedicated evaluation benchmarks for understanding negation. Most existing resources either treat negation as a minor aspect of broader tasks or focus solely on narrow syntactic detection, often emphasizing encoder-based models (hossain2020analysis; geiger2020neural; truong2022not; anschutz2023not). To address this gap, we introduce Thunder-NUBench (Negation Understanding Benchmark), a dataset explicitly designed to evaluate LLMs’ sentence-level comprehension of negation. Our benchmark is structured as a multiple-choice question (MCQ) task: given an original sentence, the model must select the correct standard negation from four options. The other three choices—local negation, contradiction, and paraphrase—are carefully designed distractors that test whether models truly grasp semantic scope and logical oppositions.
@@
  • We define standard negation within the framework of sentential logic, moving beyond the cue-based and often ambiguous accounts of prior work. Grounding standard negation in logical structure not only clarifies its role in natural language but also supports the evaluation and enhancement of reasoning in LLMs.
  • We create a manually curated benchmark that includes a dataset of sentence-negation pairs for fine-tuning, along with a multiple-choice evaluation task.
  • We conduct systematic evaluations of decoder-based LLMs, assessing their performance under both prompting and supervised fine-tuning. This includes error and confusion analyses that highlight the models’ ongoing challenges with negation.
@@ 2 Related Work
Negation detection and scope resolution.
Early work in negation detection and scope resolution primarily relied on rule-based systems and handcrafted heuristics, especially in domain-specific contexts like clinical texts. While these systems are effective, they lack flexibility across different domains (chapman2001simple; de2012ucm; ballesteros2012ucm; basile2012ugroningen). Traditional machine learning methods, such as Support Vector Machines (SVMs) (hearst1998support) and Conditional Random Fields (CRFs) (sutton2012introduction), were introduced later; however, they too are limited to narrow domains (morante2008learning; morante2009metalearning; read2012uio1; li2018learning).
More recently, deep learning approaches employing Convolutional Neural Networks (CNNs) (o2015introduction) and Bidirectional Long Short-Term Memory (BiLSTM) networks (siami2019performance) have enhanced performance by providing improved contextual embeddings and sequence modeling (fancellu2016neural; bhatia2019end). Pretrained transformer models like BERT have been employed through transfer learning techniques (e.g., NegBERT (khandelwal2020negbert)), significantly increasing the accuracy of negation detection tasks. Nonetheless, these methods still largely focus on syntactic span detection, leaving deeper semantic understanding of negation a challenging area to tackle.
@@
Negation understanding has become increasingly important in natural language understanding (NLU) tasks (hosseini2021understanding). However, existing NLU benchmarks, such as SNLI (bowman2015large) for natural language inference (NLI), CommonsenseQA (talmor2019commonsenseqa) for Question Answering (QA), SST-2 (socher2013recursive) for sentiment analysis, STS-B (cer2017semeval) for textual similarity and paraphrasing, have been criticized for not adequately addressing the semantic impact of negation (hossain2022analysis; rezaei2024paraphrasing). These datasets contain relatively few instances of negation or include negations that are not crucial to task performance, allowing language models to achieve high accuracy even when they completely ignore negation.
Recent studies, including NegNLI (hossain2020analysis), MoNLI (geiger2020neural), and NaN-NLI (truong2022not), have introduced benchmarks for NLU that are sensitive to negation. These studies show that model performance significantly declines when negation plays a crucial role in affecting the outcome (naik2018stress; yanaka2019can; hartmann2021multilingual; hossain2022question; hossain2022leveraging; she2023scone; anschutz2023not). These findings suggest that current language models tend to depend on superficial linguistic patterns rather than a genuine understanding of semantics.
@@
Distributional semantics (harris1954distributional; sahlgren2008distributional) aims to create models that learn semantic representations based on patterns of word co-occurrences (boleda2020distributional; lenci2022comparative) and capture broad semantic relationships; however, it encounters significant challenges with negation. Negated expressions, such as "not good," often appear in similar contexts as their affirmative counterparts, like "good." As a result, models tend to generate similar vector representations for these expressions, despite their opposing meanings. Previous research has pointed out this limitation, showing that PLMs struggle to capture the subtle semantic differences introduced by antonyms and the reversal of polarity (rimell2017learning; jumelet2018language; niwa2021predicting; jang2022beyond; vahtola2022not). Studies have further suggested that models like BERT find it difficult to distinguish between affirmative and negated contexts (kassner2020negated; ettinger2020bert).
@@
Recent research on understanding negation has primarily focused on bidirectional models, such as BERT (devlin2019bert) and RoBERTa (liu2019roberta), which have demonstrated strong performance in NLU and negation detection tasks. However, with the emergence of generative foundation models like GPT (radford2018improving) and LLaMA (touvron2023llama), attention has shifted towards evaluating how these models handle negation. Studies have shown that these generative models often exhibit a positive bias and struggle with producing or interpreting negated statements (truong2023language; chen2023say; garcia2023not). Although some benchmarks, such as CONDAQA (ravichander2022condaqa) and ScoNe (she2023scone), reveal these limitations, there is still a lack of robust evaluation resources specifically designed for generative models.
Building on previous studies, this paper assesses whether generative models can comprehend negation in complex sentences and identify semantic differences that extend beyond surface-level patterns.
@@ 3.1 Typology of Negation
Negation is a fundamental semantic and syntactic operation found in natural languages, used to convey denial, rejection, or the absence of a proposition. Hereafter, we denote our negation operation for a sentence $S$ as $\text{Neg}(S)$. In formal logic, negation flips the truth value of a proposition $P$: if $P$ is true, then $\text{Neg}(P)$ is false, and vice versa. Semantically, negation creates a binary opposition between a proposition and its affirmative counterpart, meaning that each one is the opposite of the other (sep-negation).
Negation can be categorized along several dimensions: scope, form, and target (see Table 1). In terms of scope, negation may affect the entire clause (referred to as clausal negation) or only part of it (known as subclausal negation). Regarding form, negation can manifest as bound morphemes, such as prefixes and suffixes (morphological negation), or as separate syntactic elements like "not" or "never" (syntactic negation). Finally, depending on its target, negation can apply to the verb (verbal negation) or to other elements in the sentence (non-verbal negation(zanuttini2001sentential; miestamo2007negation; truong2022not; kletz2023self).
@@ 3.2 Negation and Contradiction
Negation and contradiction are closely related concepts that are often conflated in NLP research (jiang2021m). Contradiction refers to the incompatibility of two propositions, meaning that they cannot both be true at the same time. While negation frequently serves as a primary mechanism for creating contradictions—by reversing the truth value of a proposition—contradictions can also arise from antonymy, numeric mismatches, or differences in structure and lexicon (further details can be found in Appendix A). For instance, the statements "An individual was born in France" and "An individual was born in Italy" are contradictory, but they are not negations, as the second statement does not reverse the truth of the first.
@@
The man who owns the car
is my neighbor. $\rightarrow$ The man
who does not own the car is
my neighbor.
@@
Walking through the park, she
found a lost wallet. $\rightarrow$ Not
walking through the park, she
@@
raining. $\rightarrow$ She stayed inside
because it was not raining.
@@
attended the meeting. $\rightarrow$ He
submitted the report and did
@@ 3.3 Standard Negation
Standard negation refers to the typical form of negation applied to the declarative verbal main clause. It specifically negates the verb in a main clause (miestamo2000towards). A main clause can function as a complete sentence on its own, consisting at minimum of a subject and a predicate. This definition is grounded in the notion that the verb acts as the head of the clause (miller2011critical).
@@
Building on this traditional understanding, we treat standard negation as the process of reversing the truth value of the verb phrase in the main clause, which we will refer to as the main predicate in this paper. A verb phrase is headed by a verb and can consist of a single verb or a combination of auxiliaries, complements, and modifiers (e.g., "will call" and "is being promoted") (lakoff1966criterion). Since the main predicate conveys the core action or state of the clause, negating it effectively reverses the proposition of the entire sentence. In this context, we treat standard negation as a truth-functional operation that maps the main predicate to its complement set within the semantic space.
We further clarify the scope of standard negation within the typology presented in Table 1. Standard negation includes both clausal negation and verbal negation, as it reverses the meaning of the entire sentence by negating the main predicate. In terms of form, standard negation can employ both syntactic and morphological negation.
Syntactically, standard negation often uses explicit negation particles, such as "not." Morphologically, it can involve complementary antonyms (for example, "alive" vs. "dead" or "true" vs. "false"), which occupy mutually exclusive semantic spaces, thus reversing the truth value of the proposition.
In contrast, other types of antonyms, such as gradable antonyms (e.g., "happy" vs. "unhappy") and relational antonyms (e.g., "buy" vs. "sell") (lehrer1982antonymy), do not strictly reverse truth values. Therefore, they are classified as contradictions rather than standard negation in this paper.
Atomic propositions.
While this characterization effectively defines standard negation for atomic propositions (elementary sentences that cannot be further decomposed) (davis2004semantics), its application to complex sentences with multiple clauses requires a more thorough approach. In this paper, we treat an atomic proposition as a sentence that contains a single main predicate.
Complex propositions.
@@
Specifically, for propositions composed of multiple logically connected atomic statements, the method for reversing the truth value of the entire complex proposition can be ambiguous. In natural language, such logical structures typically appear as coordinated clauses (e.g., "$P$ and $Q$ or RR") or comma-separated lists connected by "and" or "or" (e.g., "$P,Q,$ and RR"). We treat these as equivalent to a sequence of binary conjunctions or disjunctions.
@@
Definition of standard negation.
In this paper, standard negation refers to natural-language sentential negation, which is formally treated as logical negation within the framework of sentential logic (enderton2001). To address the complexities involved, we define standard negation recursively by applying it pairwise over the logical structure of a sentence until only atomic propositions remain, ensuring that the truth value of the entire sentence is reversed even when it contains multiple coordinated clauses. Our definition of $\text{Neg}(\cdot)$ is presented in Table 2. Conditionals of the form "if $P$, QQ" are equivalent to $\text{Neg}(P)\;\text{or}\;Q$" in logic, and we adhere to this equivalence when defining their negation (more details can be found in Appendix C).
Our definition of standard negation is inspired by the framework of sentential logic (enderton2001). However, it should be viewed as an operational definition rather than a strict mathematical formulation. Natural language sentences often lack explicit structural markers, such as parentheses, which are vital in well-formed logical formulas. Furthermore, coordination can appear with or without commas. Unlike formal languages, natural languages do not adhere to strict formation rules, making it challenging to map their structures unambiguously to logical formulas. Consequently, our definition cannot perfectly align with the forms used in sentential logic. Nevertheless, it provides a clear operational account of standard negation in natural languages.

### v4

@@ Abstract
Negation is a fundamental linguistic phenomenon that poses ongoing challenges for Large Language Models (LLMs), particularly in tasks requiring deep semantic understanding. Current benchmarks often treat negation as a minor detail within broader tasks, such as natural language inference. Consequently, there is a lack of benchmarks specifically designed to evaluate comprehension of negation. In this work, we introduce Thunder-NUBench — a novel benchmark explicitly created to assess sentence-level understanding of negation in LLMs. Thunder-NUBench goes beyond identifying surface-level cues by contrasting standard negation with structurally diverse alternatives, such as local negation, contradiction, and paraphrase. This benchmark includes manually created sentence-negation pairs and a multiple-choice dataset, allowing for a comprehensive evaluation of models’ understanding of negation.
Thunder-NUBench: A Benchmark for LLMs’
Sentence-Level Negation Understanding
Yeonkyoung So1, Gyuseong Lee1, Sungmok Jung1, Joonhak Lee1, JiA Kang1, Sangho Kim1, Jaejin Lee1,2 1Graduate School of Data Science, Seoul National University 2Dept. of Computer Science and Engineering, Seoul National University {kathy1028,ksnannaya,tjdahrwjd,hmjelee,jia6776,ksh4931,jaejin}@snu.ac.kr
@@ 1 Introduction
Negation is a fundamental and universal phenomenon found in languages worldwide. It is closely associated with various human communicative abilities, such as denial, contradiction, deception, misrepresentation, and irony. Although affirmative statements are more common, negation still plays a significant role in language; approximately 25% of sentences in English texts contain some form of negation (Sarabi and Blanco, 2016; Hossain et al., 2020; Horn and Wansing, 2025). This prevalence and its impact on meaning make accurate interpretation of negation crucial for several natural language processing (NLP) tasks, including sentiment analysis, question answering, knowledge base completion, and natural language inference (NLI) (Khandelwal and Sawant, 2020; Hosseini et al., 2021; Singh et al., 2023). Recent studies have shown that effectively managing negation is important even for multi-modal language models (Quantmeyer et al., 2024; Alhamoud et al., 2025; Park et al., 2025).
Figure 1: Illustration of the Thunder-NUBench task. Only the standard negation reverses the truth value of the original sentence, while other options differ in scope or semantics.
Meanwhile, negation poses significant challenges for both humans and language models. Research shows that people often find it more difficult to process and comprehend negated statements compared to affirmative ones (Wales and Grieve, 1969; Sarabi and Blanco, 2016). Similarly, many studies indicate that pretrained language models (PLMs) struggle to interpret negation accurately. For example, models like BERT (Devlin et al., 2019) and even large language models (LLMs) such as GPT-3 (Brown et al., 2020) often have difficulty distinguishing between negated and affirmative statements. These models tend to rely on superficial cues, which can result in incorrect outputs when negation is involved (Kassner and Schütze, 2020; Hossain et al., 2022a; Truong et al., 2023).
Despite its significance, there is a notable lack of dedicated evaluation benchmarks for understanding negation. Most existing resources either treat negation as a minor aspect of broader tasks or focus solely on narrow syntactic detection, often emphasizing encoder-based models (Hossain et al., 2020; Geiger et al., 2020; Truong et al., 2022; Anschütz et al., 2023). To address this gap, we introduce Thunder-NUBench (Negation Understanding Benchmark), a dataset explicitly designed to evaluate LLMs’ sentence-level comprehension of negation. [^1] Thunder-NUBench is publicly available at https://huggingface.co/datasets/thunder-research-group/SNU_Thunder-NUBench. Our benchmark is structured as a multiple-choice question (MCQ) task: given an original sentence, the model must select the correct standard negation from four options. The other three choices (local negation, contradiction, and paraphrase) are carefully designed distractors that test whether models truly grasp semantic scope and logical oppositions.
@@
  • We define standard negation within the framework of sentential logic. Grounding standard negation in logical structure not only clarifies its role in natural language but also supports the evaluation and enhancement of reasoning in LLMs.
  • We introduce a manually created benchmark that includes a dataset of sentence-negation pairs for fine-tuning, along with a multiple-choice evaluation task.
  • We conduct systematic evaluations of decoder-based LLMs across model families, scales, and training methods to analyze variation in negation understanding.
@@ 2 Related Work
Negation detection and scope resolution.
Early work in negation detection and scope resolution primarily relied on rule-based systems and handcrafted heuristics, especially in domain-specific contexts like clinical texts. While these systems are effective, they lack flexibility across different domains (Chapman et al., 2001; de Albornoz et al., 2012; Ballesteros et al., 2012; Basile et al., 2012). Traditional machine learning methods, such as Support Vector Machines (SVMs) (Hearst et al., 1998) and Conditional Random Fields (CRFs) (Sutton et al., 2012), were introduced later; however, they too are limited to narrow domains (Morante et al., 2008; Morante and Daelemans, 2009; Read et al., 2012; Li and Lu, 2018).
More recently, deep learning approaches employing Convolutional Neural Networks (CNNs) (O’shea and Nash, 2015) and Bidirectional Long Short-Term Memory (BiLSTM) networks (Siami-Namini et al., 2019) have enhanced performance by providing improved contextual embeddings and sequence modeling (Fancellu et al., 2016; Bhatia et al., 2019). Pretrained transformer models like BERT have been employed through transfer learning techniques (e.g., NegBERT (Khandelwal and Sawant, 2020)), significantly increasing the accuracy of negation detection tasks. Nonetheless, these methods still largely focus on syntactic span detection, leaving deeper semantic understanding of negation a challenging area to tackle.
@@
Negation understanding has become increasingly important in natural language understanding (NLU) tasks (Hosseini et al., 2021). However, existing NLU benchmarks, such as SNLI (Bowman et al., 2015) for natural language inference (NLI), CommonsenseQA (Talmor et al., 2019) for Question Answering (QA), SST-2 (Socher et al., 2013) for sentiment analysis, STS-B (Cer et al., 2017) for textual similarity and paraphrasing, have been criticized for not adequately addressing the semantic impact of negation (Hossain et al., 2022a; Rezaei and Blanco, 2024). These datasets contain relatively few instances of negation or include negations that are not crucial to task performance, allowing language models to achieve high accuracy even when they completely ignore negation.
Recent studies, including NegNLI (Hossain et al., 2020), MoNLI (Geiger et al., 2020), NaN-NLI (Truong et al., 2022), NoFEVER-ML and NoSNLI-ML (Vrabcová et al., 2025), have introduced benchmarks for NLU that includes negation. These studies show that model performance significantly declines when negation plays a crucial role in affecting the outcome (Naik et al., 2018; Yanaka et al., 2019; Hartmann et al., 2021; Hossain et al., 2022b; Hossain and Blanco, 2022; She et al., 2023; Anschütz et al., 2023). These findings suggest that current language models tend to depend on superficial linguistic patterns rather than a genuine understanding of semantics.
@@
Distributional semantics (Harris, 1954; Sahlgren, 2008) aims to create models that learn semantic representations based on patterns of word co-occurrences (Boleda, 2020; Lenci et al., 2022) and capture broad semantic relationships; however, it encounters significant challenges with negation. Negated expressions, such as "not good," often appear in similar contexts as their affirmative counterparts, like "good." As a result, models tend to generate similar vector representations for these expressions, despite their opposing meanings. Previous research has pointed out this limitation, showing that PLMs struggle to capture the subtle semantic differences introduced by antonyms and the reversal of polarity (Rimell et al., 2017; Jumelet and Hupkes, 2018; Niwa et al., 2021; Jang et al., 2022; Vahtola et al., 2022). Studies have further suggested that models like BERT find it difficult to distinguish between affirmative and negated contexts (Kassner and Schütze, 2020; Ettinger, 2020).
@@
Recent research on understanding negation has primarily focused on bidirectional models, such as BERT (Devlin et al., 2019) and RoBERTa (Liu et al., 2019), which have demonstrated strong performance in NLU and negation detection tasks. However, with the emergence of generative foundation models like GPT (Radford et al., 2018) and LLaMA (Touvron et al., 2023), attention has shifted towards evaluating how these models handle negation. Studies have shown that these generative models often exhibit a positive bias and struggle with producing or interpreting negated statements (Truong et al., 2023; Chen et al., 2023; García-Ferrero et al., 2023). Although some benchmarks, such as CONDAQA (Ravichander et al., 2022) and ScoNe (She et al., 2023), reveal these limitations, there is still a lack of robust evaluation resources specifically designed for negation understanding of generative models.
Building on previous studies, this paper focuses on sentence-level negation as a core linguistic operation. While prior negation-related NLI or QA benchmarks evaluate negation within broader inference or comprehension tasks, Thunder-NUBench specifically tests whether models can distinguish standard negation, the logical reversal of the sentence, from closely related distractors. As a result, Thunder-NUBench enables a more direct examination of sentence-level negation than is afforded by NLI or QA benchmarks, where negation is evaluated only as part of broader inference tasks.
@@ 3.1 Typology of Negation
Negation is a fundamental semantic and syntactic operation found in natural languages, used to convey denial, rejection, or the absence of a proposition. Hereafter, we denote our negation operation for a sentence $S$ as $\text{Neg}(S)$. In formal logic, negation flips the truth value of a proposition $P$: if $P$ is true, then $\text{Neg}(P)$ is false, and vice versa. Semantically, negation creates a binary opposition between a proposition and its affirmative counterpart, meaning that each one is the opposite of the other (Horn and Wansing, 2025).
Negation can be categorized along several dimensions: scope, form, and target (see Table 1). In terms of scope, negation may affect the entire clause (referred to as clausal negation) or only part of it (known as subclausal negation). Regarding form, negation can manifest as bound morphemes, such as prefixes and suffixes (morphological negation), or as separate syntactic elements like "not" or "never" (syntactic negation). Finally, depending on its target, negation can apply to the verb (verbal negation) or to other elements in the sentence (non-verbal negation(Zanuttini, 2001; Miestamo, 2007; Truong et al., 2022; Kletz et al., 2023).
@@ 3.2 Negation and Contradiction
Negation and contradiction are closely related concepts that are often conflated in NLP research (Jiang et al., 2021). Contradiction refers to the incompatibility of two propositions, meaning that they cannot both be true at the same time. While negation frequently serves as a primary mechanism for creating contradictions (by reversing the truth value of a proposition), contradictions can also arise from antonymy, numeric mismatches, or differences in structure and lexicon (further details can be found in Appendix A). For instance, the statements "An individual was born in France" and "An individual was born in Italy" are contradictory, but they are not negations, as the second statement does not reverse the truth of the first.
@@
The man who owns the car is my
    neighbor.
$\rightarrow$ The man who does not own the car
     is my neighbor.
@@
Walking through the park, she found
    a lost wallet.
$\rightarrow$ Not walking through the park, she
@@
    raining.
$\rightarrow$ She stayed inside because it was
    not raining.
@@
    attended the meeting.
$\rightarrow$ He submitted the report and did
@@ 3.3 Standard Negation
Standard negation refers to the typical form of negation applied to the declarative verbal main clause. It specifically negates the verb in a main clause (Miestamo, 2000). A main clause can function as a complete sentence on its own, consisting at a minimum of a subject and a predicate. This definition is grounded in the notion that the verb acts as the head of the clause (Miller and Miller, 2011).
@@
Building on this traditional understanding, we treat standard negation as the process of reversing the truth value of the verb phrase in the main clause, which we will refer to as the main predicate in this paper. A verb phrase is headed by a verb and can consist of a single verb or a combination of auxiliaries, complements, and modifiers (e.g., "will call" and "is being promoted") (Lakoff, 1966). Since the main predicate conveys the core action or state of the clause, negating it effectively reverses the proposition of the entire sentence. In this context, this paper treats standard negation as a truth-functional operation that maps the main predicate to its complement set within the semantic space.
We further clarify the scope of standard negation within the typology presented in Table 1. Standard negation includes both clausal negation and verbal negation, as it reverses the meaning of the entire sentence by negating the main predicate. In terms of form, standard negation can employ both syntactic and morphological negation. Syntactically, standard negation often uses explicit negation particles, such as "not." Morphologically, it can involve complementary antonyms (for example, "alive" vs. "dead" or "true" vs. "false"), which occupy mutually exclusive semantic spaces, thus reversing the truth value of the proposition. In contrast, other types of antonyms, such as gradable antonyms (e.g., "happy" vs. "unhappy") and relational antonyms (e.g., "buy" vs. "sell") (Lehrer and Lehrer, 1982), do not strictly reverse truth values. Therefore, they are classified as contradictions rather than standard negation in this paper.
Atomic and Complex Propositions.
@@
While this characterization of standard negation effectively defines standard negation for atomic propositions (elementary sentences that cannot be further decomposed) (Davis and Gillon, 2004), its application to complex sentences with multiple clauses requires a more thorough approach. In this paper, we treat an atomic proposition as a sentence that contains a single main predicate. Specifically, for propositions composed of multiple logically connected atomic statements, the method for reversing the truth value of the entire complex proposition can be ambiguous. In natural language, such logical structures typically appear as coordinated clauses (e.g., "$P$ and $Q$ or RR") or comma-separated lists connected by "and" or "or" (e.g., "$P,Q,$ and RR"). We treat these as equivalent to a sequence of binary conjunctions or disjunctions.
Definition of standard negation.
In this paper, standard negation refers to natural-language sentential negation, which is formally treated as logical negation within the framework of sentential logic (Enderton, 2001). To address the complexities involved, we define standard negation recursively by applying it pairwise over the logical structure of a sentence until only atomic propositions remain, ensuring that the truth value of the entire sentence is reversed even when it contains multiple coordinated clauses. Our definition of $\text{Neg}(\cdot)$ is presented in Table 2. Conditionals of the form "if $P$, QQ" are equivalent to $\text{Neg}(P)\;\text{or}\;Q$" in logic, and we adhere to this equivalence when defining their negation (more details can be found in Appendix C).